How long must a guywire be to run from the top of an 18-ft pole to a point on level ground 16 ft away from the base of the pole?
Answer by checkley71(8403)

HERE YOU HAVE A RIGHT TRIANGLE WITH SIDES OF 18 7 16 FEET. THE GUYWIRE IS THE HYPOTENUSE.
USING THE PYTHAGOREON FORMULA WE GET:
18^2+16^2=H^2
324+256=H^2
580=H^2
H=SQRT580
H=24.08 FEET FOR THE GUYWIRE.